### v2

# Judging from Support-set: A New Way to Utilize Few-Shot Segmentation for Segmentation Refinement Process

Abstract
Segmentation refinement aims to enhance the initial coarse masks generated by segmentation algorithms. The refined masks are expected to capture more details and better contours of the target objects. Research on segmentation refinement has developed as a response to the need for high-quality image segmentations. However, to our knowledge, no method has been developed that can determine the success of segmentation refinement. Such a method could ensure the reliability of segmentation in applications where the outcome of the segmentation is important and fosters innovation in image processing technologies. To address this research gap, we propose Judging From Support-set (JFS), a method to judge the success of segmentation refinement leveraging an off-the-shelf few-shot segmentation (FSS) model. The traditional goal of the problem in FSS is to find a target object in a query image utilizing target information given by a support set. However, we propose a novel application of the FSS model in our evaluation pipeline for segmentation refinement methods. Given a coarse mask as input, segmentation refinement methods produce a refined mask; these two masks become new support masks for the FSS model. The existing support mask then serves as the test set for the FSS model to evaluate the quality of the refined segmentation by the segmentation refinement methods.We demonstrate the effectiveness of our proposed JFS framework by evaluating the SAM Enhanced Pseduo-Labels (SEPL) using SegGPT as the choice of FSS model on the PASCAL dataset. The results showed that JFS has the potential to determine whether the segmentation refinement process is successful.
Figure 1: (a) The goal of conventional few-shot segmentation is to find the target object in the query image when a support set is given. (b) JFS approach. When there are a coarse mask and a refined mask, we can use the FSS network and support-set to prove that the refinement process is successful. Query image, segmentation masks (Coarse mask prediction and refined prediction), and support-set used for deciding the success of refinement segmentation. Note: In this figure, the error analysis for a refined mask is only visualized.
Figure 2: Segmentation refinement is a critical stage that aims to improve the initial, coarse masks generated by segmentation algorithms. These refined masks are supposed to capture better details and contours of the target objects. In the segmentation refinement process, as shown in the Stage 1 result on the left, problems can occur when the refined mask results in post-processing fails. The existing segmentation refinement has major limitations in its inability to determine whether refinement has failed because of non-existing ground truth information. We propose Judging from Support Set (JFS), which can determine whether the refined mask has better results than the previous coarse mask or not. In Stage 2, Support mask prediction-C ($SMP^{c}$) is the result obtained using a coarse mask, and support mask prediction-R ($SMP^{r})$ is the result obtained using a refined mask. The cases in Stage 2 show when the support mask prediction from the coarse mask performs better than the refined mask. $SMP^{c}$ closely matches the support mask. JFS can identify these cases by comparing the performance of coarse mask and refined mask in Stage 1 using FSS network. Therefore, we can identify whether segmentation refinement is successful or failed.
Introduction
Semantic segmentation, which involves labeling each pixel of an image with a semantic category, is widely used in numerous fields, including medical imaging (Yao et al. 2023), semiconductor manufacturing (Huang and Pan 2015), and autonomous driving (Muhammad et al. 2022). Recent advancements in Deep Convolutional Neural Networks (DCNNs) and the availability of extensive pixel-level annotated data have significantly enhanced the effectiveness of semantic segmentation techniques (Chen et al. 2018). However, due to the complexity of directly generating high-quality masks, some prior research has focused on improving coarse masks generated by initial segmentation models. Previous works (Kirillov et al. 2020; Ke et al. 2021) involve improving segmentation models with a custom module designed for mask correction. Other segmentation refinement studies (Yuan et al. 2020; Chen et al. 2023) adopt model-agnostic approaches, using only an image and its coarse segmentation as input for refinement. Likewise, segmentation refinement continues to develop through various methods.
Indeed, segmentation refinement can improve most initial segmentations; however, it is important to note that segmentation refinement is not guaranteed to be successful (Figure 2) and the best results can only be expected when selecting an appropriate refinement model that performs well on a specific task. This makes determining the success and failure of segmentation refinement important. If the success and failure of segmentation refinement can be monitored, it would help to identify the most effective refinement techniques and achieve the best semantic segmentation. However, in contrast to the rapid developments of segmentation refinement techniques, we point out that the method for identifying whether the refinement is successful has been understudied despite its importance and usefulness. To address this research gap, we propose a new method that can assess the success of segmentation refinement using a few-shot segmentation (FSS) network.
FSS is a class of methods that aims to address the shortcomings of supervised learning segmentation models. Specifically, FSS segments the target object in a query image using a few support image-segmentation pairs as prompts to reduce the cost and effort of curating new sets of data for retraining segmentation models from scratch. We found that FSS can also be applied to evaluate segmentation refinement methods. Since the support set prompting the FSS model contains ground-truth segmentation already, we can use them as queries instead and prompting the FSS model with the refined mask from the segmentation refinement modules. By comparing and evaluating the output of the FSS model prompted by the refined mask and the original coarse mask, we can determine whether the refined mask is truly an improvement over the coarse mask. Based on this idea, we proposed a new method that can determine whether the segmentation refinement process is successful by using the support set and query images inversely from the conventional FSS approach (Figure 1). This study contributes to the field of segmentation refinement by offering a novel way to analyze the quality of refinement masks.
To the best of our knowledge, we are the first to introduce a method of using the FSS model for analyzing the segmentation refinement process. Our method is named JFS (Judging From Support-Set) in the sense that it uses information from the support set in the FSS to determine the success of the refinement. Our proposed method can determine whether segmentation refinement was successful or failed. We summarize our key contributions as follows:
  • We proposed JFS, which is able to leverages an off-the-shelf FSS network to evaluate segmentation refinement. We believe this is the first method to determine the success of segmentation refinement methods.
  • We verify the effectiveness of our proposed method by combining JFS with the SAM Enhanced pseudo-label (SEPL) pipeline on worst and best cases in PASCAL (Everingham et al. 2010).
Related work
Segmentation Refinement
Segmentation refinement aims to improve the quality of masks generated by existing segmentation models. There are two main approaches to segmentation refinement: model-specific methods and model-agnostic methods. Model-specific methods involve designing algorithms that are tailored to a specific type of model. For instance, PointRend (Kirillov et al. 2020) and MaskTransfiner (Ke et al. 2021) are examples of model-specific methods. PointRend (Kirillov et al. 2020) conducts point-based segmentation predictions at dynamically chosen locations using an iterative subdivision algorithm. MaskTransfiner (Ke et al. 2021) constructs a quadtree from the sparse and incoherent regions within the RoI pyramid, and subsequently refines all tree nodes collectively using a refinement transformer that incorporates quadtree attention. Unlike model-specific methods, model-agnostic methods are more versatile in that they can refine masks from various models. For example, SegFix (Yuan et al. 2020) utilized boundary loss and direction loss separately on the predicted boundary map and direction map, respectively. BPR (Tang et al. 2021) extracts and enhances a series of boundary patches along the predicted instance boundaries using a powerful refinement network. SEPL (Chen et al. 2023) employs CAM pseudo-labels as indicators to identify and merge SAM masks, producing high-quality pseudo-labels that are aware of both class and object. Segmentation refinement techniques have been widely used to improve coarse masks. However, no method has been developed that can identify the effectiveness of post-processing in improving initial segmentation. This study aims to address this research gap by proposing a new method, JFS.
Few-shot Segmentation
Few-shot segmentation (FSS) aims to accurately segment a target object in a query image using a limited number of annotated examples called support-set. FSS was introduced by Shaban et al. (Shaban et al. 2017). Since then, many different methods have been attempted to improve FSS performance. The CANet (Zhang et al. 2019) enhanced performance through iterations and cosine similarity adjustments, while PFENet (Tian et al. 2020) and PANet (Wang et al. 2019) focused on improving segmentation quality with pyramid modules and prototype networks, respectively. The other methods such as PPNet (Liu et al. 2020) and PMM (Yang et al. 2020) leveraged detailed object features and multiple prototypes for accuracy. ASGNet (Li et al. 2021) introduced modules for better prototype extraction and allocation, and DAN (Wang et al. 2020) focused on improving support and query image correspondence. HSNet (Min, Kang, and Cho 2021) proposed a 4D convolution for feature analysis, and VAT (Hong et al. 2022) proposed a transformer-based network to model dense semantic correspondence. Various studies are continuously being conducted to improve FSS accuracy, but to our knowledge, there is no previous research using FSS to determine whether segment refinement is successful.
Methodology
The overall process of JFS is shown in Figure 3. First, we locate the target object using a segmentation model to obtain a coarse mask prediction. At this stage, the segmentation mask is often a coarse one. This is followed by the SAM Enhanced Pseduo-Labels (SEPL) (Chen et al. 2023) refinement process. During the refinement, the use of a robust Segmentation model, SAM (Kirillov et al. 2023), allows for the acquisition of a more accurate segmentation mask of the object. However, whether the refined result is more accurate than the original prediction is questionable. The final step of our pipeline aims to answer this question. Specifically, we propose utilizing the FSS model to determine whether the refined segmentation is successful. The proposed framework is divided into two stages: (1) Coarse Mask Generation and Refinement and (2) Judging from Support set. This initial stage focuses on generating a preliminary segmentation mask, termed the ‘coarse mask’. Subsequently, this mask undergoes a refinement process to enhance its accuracy and detail, resulting in a ‘refined mask’. The second stage involves evaluating the refined masks by using them as support sets for FSS model. In this stage, another set of images along with their ground-truth segmentation is required for the evaluation, and we choose the original support set for the FSS model. These stages will be elaborated upon in the following discussion.
Figure 3: Overview of JFS. JFS uses support mask $M^{s}$ as ground truth to determine which mask prediction is closer to the ground truth. If the IoU value of $E_{c}$ is greater than $E_{r}$, segmentation refinement has failed. Conversely, if the IoU value of $E_{r}$ is greater than $E_{c}$, segmentation refinement has succeeded. Note: The terms Query role and Support-set role refer to their respective roles in Few-Shot Segmentation.
Stage 1: Coarse Mask Generation and Refinement
We utilized CAM (Li et al. 2023) to generate a coarse mask and the generated coarse mask is refined using SEPL (Chen et al. 2023). Coarse masks generated by CAM recognize distinct, class-specific regions but often miss precise object boundaries. We do not expect an accurate coarse mask because the purpose of the proposed method, JFS, is to determine whether segment refinement is successful or not. To get a refined mask, we utilized SEPL which improves the quality of coarse masks by using a strong segmentation tool, the Segment Anything Model (SAM) (Kirillov et al. 2023). Because SAM can accurately segment most parts or objects regardless of their class, SEPL can produce a high-quality segmentation mask. For one image, SEPL takes input coarse mask $[M^{c}_{1},\cdots,M^{c}_{K}]$ and SAM masks $[S_{0},\cdots,S_{L}]$ and outputs refined masks $[M^{r}_{1},\cdots,M^{r}_{K}]$. This process involves two steps: mask assignment and mask selection. In the mask assignment stage, SEPL calculates the intersection of each SAM mask $S_{l}$ with the coarse mask $M^{c}_{k}$ for each class $k$ from $1$ to $k$. Each mask $S_{l}$ is assigned to the class with which it overlaps the most, while masks with no overlap are discarded.
Note that although SEPL (Chen et al. 2023) generally improves segmentation quality, it has drawbacks from three sources: when a coarse mask fails to activate a target object, when the SAM fails to detect a target object, and the algorithm itself tends to have a large segment. These fail to improve segmentation quality from coarse masks and SEPL itself does not provide any information if refined masks have a worse quality mask than coarse masks.
Stage 2: Judging from Support Set
In this stage, we determine whether the refinement process is successful or not using the Algorithm 1 (Figure 3). With the two masks generated from stage 1, the coarse mask $M^{c}$ and the refined mask $M^{r}$, the query image $I^{q}$ (Original image used to create the coarse mask), the support image $I^{s}$ and the support mask $M^{s}$ are input to JFS. We avoid using the same image from PASCAL (Everingham et al. 2010) for query and support images. Therefore, the query images are from the PASCAL train set, and the support image $I^{s}$ was taken from the PASCAL validation set. The support image should contain the same class object in the query image. All these inputs go to SegGPT (Wang et al. 2023) which has shown a strong performance in the FSS task to get accurate support mask prediction ($SMP$). We get $SMP^{c}$ and $SMP^{r}$ from $M^{c}$ and $M^{r}$, respectively. We can obtain the IoU value by comparing the results of $SMP$ and $M^{s}$ using Equation 1. The higher the IoU value, the more similar it is to $M^{s}$. When $E_{c}$ is greater than $E_{r}$, it indicates that the segmentation refinement has failed. In contrast, if the value of $E_{r}$ is greater than $E_{c}$, it suggests that the refinement was successful.
$E_{c}=\frac{SMP^{c}\cap M^{s}}{SMP^{c}\cup M^{s}},\quad E_{r}=\frac{SMP^{r}\cap M^{s}}{SMP^{r}\cup M^{s}}$ (1)
Note that other few-shot segmentation models can be utilized for this stage not only SegGPT.
Algorithm 1 JFS: Judging from Support Set
Input: Query Image $I^{q}$, Coarse Mask $M^{c}$, Refined Mask $M^{r}$, Support Image $I^{s}$, Support Mask $M^{s}$
Output: $E_{c}$ and $E_{r}$ scores
1:  Generate two sets of data A and B. $A=\{I^{q},M^{c},I^{s}\}$, $B=\{I^{q},M^{r},I^{s}\}$
2:  Both $I^{q},M^{c}$ and $I^{q},M^{r}$ behave as Support-set role and $I^{s}$ becomes a query image in FSS
3:  Get Support Mask Prediction $SMP^{c}$ from A and $SMP^{r}$ from B using FSS network (SegGPT)
4:  Calculate $E_{c}$ and $E_{r}$ comparing $SMP^{c}$, $SMP^{r}$ with $M^{s}$ respectively using Equation 1
Experiment
Design: To evaluate JFS’s ability to judge refined segmentations, we designed the following experiment. We use JFS to determine whether a refined mask improves over a coarse mask. If so, the refined mask is kept; otherwise, the coarse mask is retained. This process should result in a set of segment masks of higher quality by combining the best among the coarse and refined cases. We then measure the IOU of this new set against the original coarse and refined masks to assess the quality improvement achieved by JFS’s judgment. We exclude samples where both the coarse mask and refined mask have an IoU of 0.
Datasets: For quantatative analysis, we evaluated the JFS method on the PASCAL (Everingham et al. 2010) dataset. We generate coarse $M^{c}$ and refined mask $M^{r}$ for images $I^{q}$ in the PASCAL training set; then generate few-shot segmentation prediction $SMP^{c}$ and $SMP^{r}$ for the validation set images $I^{s}$ using Algorithm 1 . We perform analysis on several groups of instances: 1) Randomly select 400 images, 20 for each of the 20 classes; 2) Top and Bottom 50 improved SEPL outputs; 3) Top 20 improved SEPL outputs and 4) Bottom 20 improved SEPL outputs. [^1] Improved means IoU difference between SEPL refined and original coarse segmentations In addition, we use gathered license-free images per class for support-set from Wikimedia Commons for qualitative results.
Evaluation metrics: We compared support mask prediction performance with GT support mask using Intersection-over-Union (IoU) which is widely used in segmentation tasks (Min, Kang, and Cho 2021; Kang and Cho 2022; Hong et al. 2022; Moon et al. 2022). IoU calculates the overlap between the predicted region (mask) and the ground truth region (mask) of an object in an image, providing a measure of how accurately the predicted mask aligns with the ground truth mask.
$\text{IoU}=\frac{|A\cap B|}{|A\cup B|}$ (2)
$A$ represents the ground truth mask. $B$ represents the prediction mask. $|A\cap B|$ is the intersection area between the ground truth and prediction masks. $|A\cup B|$ is the area of union of the ground truth and prediction masks. We calculate mean IOU, $mIoU=\frac{1}{n}\sum_{1}^{n}IoU$ where $n$ is the number of test cases.
Implementation Details: Coarse masks were created using TransCAM (Li et al. 2023). We use SEPL (Chen et al. 2023) method for refinement. All SAM (Kirillov et al. 2023) masks used in SEPL (Chen et al. 2023) were generated using the pre-trained VIT-H model ((Dosovitskiy et al. 2021)) shared on the SAM official site. SegGPT was used for the FSS model, and the pre-trained model (VIT large) was used without fine-tuning.
Figure 4: This shows how JFS works when segmentation refinement is successful. Support Mask Prediction-R ($SMP^{r}$) from the refined mask, which is closer to the ground truth’s support mask $M^{s}$, is the result of more accurate segmentation of the target object in Stage 1 using SegGPT. Therefore, JFS will determine that segmentation refinement was successful in these cases.
Result and Discussion
Data mIoU mIoU (SEPL) mIoU (JFS) Success Rate
[…]
Table 1: Comparison of the mIoU of the coarse mask, the SEPL refined mask, and JFS picked segmentations, along with the success rate of JFS across different data selections. The success rate was measured as the number of correct decisions out of the total samples. Top 50/20 refers to the 50/20 samples with the greatest improvement in the refinement process, while Low 50/20 denotes the 50/20 samples with the least improvement. The best results are highlighted in bold.
Quantitative results: Table 1 compares the mIoU for coarse, refine and JFS picked segmentations, and shows the success rate of JFS across different experiment groups. The results show that the JFS picked segmentations consistently outperform the SEPL refinement in all four groups. The highest improvement by JFS was observed in the low 20 samples (where SEPL struggles), where JFS improved mIoU to 50.5 from SEPL’s 33.2. In this group, both methods showed lower mIoU than the original coarse masks, but JFS improves on SEPL outputs by detecting failure cases in SEPL outputs. Furthermore, the success rate, which measures the effectiveness of JFS, is higher where mIoU is significantly improved by SEPL.
Qualitative results: Figure 2 and Figure 4 show that JFS can determine failure and success cases in the segmentation refinement process. Therefore, if the segmentation refinement fails, we can return to the existing coarse mask and consider other segmentation refinement methods. For this visualization, we used images publicly available on Wikimedia Commons for support-set. Query images are from PASCAL. We selected images where only one object was visible for each class. The support mask was obtained using the SAM segmentation method.
How JFS works?: Most FSS networks (Shaban et al. 2017; Zhang et al. 2019; Tian et al. 2020; Liu et al. 2020; Wang et al. 2020; Min, Kang, and Cho 2021; Hong et al. 2022; Moon et al. 2022, 2023) rely on high quality support set information to find target objects in the query image (Figure 1). Therefore, FSS performance deteriorates when the support mask information is inaccurate (Moon et al. 2022, 2023). If a refined mask accurately segments the target object, it provides more detailed texture information of the target to FSS model. Hence, we designed JFS. We found from our qualitative analysis (Figure 2 and Figure 4) that the JFS method was particularly effective in the following cases :
(a) When refined masks accurately segment the target object boundaries, support mask predictions will perform better at distinguishing the target object from the background.
(b) When the difference between the coarse and refined mask is clear, the presence of non-target objects in the refined mask confuses the FSS model, significantly dropping support mask prediction accuracy. Thus, it becomes evident whether the refined mask has failed or succeeded. Table 1 shows the JFS success rate rises with significant mIoU improvement, indicating JFS’s higher effectiveness when refinement module makes greater improvement.
(c) When the query image and support-set have the same class object in the background. If the coarse mask or refined mask segments the target object and an object in the background, FSS models recognize this information as target information and attempt to segment it together if a similar object exists in the support image.
Limitations of JFS: Table 1 suggests that the success rate of JFS decreases when there is a minor improvement in the refinement process. Specifically, our observations reveal the limitations of JFS in two cases (Figure 5).
(a) If the refined mask is not significantly different from the existing segmentation, JFS struggles as $E^{c}$ and $E^{r}$ values are similar. Therefore, JFS performs better in deciding on cases where the refined mask shows a significant difference from the coarse mask.
(b) When both the refined mask and the coarse mask share the key feature of the object, FSS network will generate similar support mask predictions and the results of JFS need to be carefully considered in such cases.
Figure 5: Failure cases. When the refined mask and coarse mask cover the same key features on the same target object, support mask prediction yields similar results. This is because FSS models are trained to find a target even with small key features from support-set. JFS failed to decide accurately whether the refinement process was successful or failed.
Path to Deployment
Addressing the limitations outlined above is crucial to improving the performance of JFS. There are two potential strategies to improve the performance of JFS. The first approach is to develop and integrate a more effective segmentation refinement tool. A robust refinement tool could significantly enhance the precision and reliability of the initial segmentation output, laying a stronger foundation for subsequent processes. The second approach is to deploy a more accurate few-shot segmentation (FSS) model in the second stage of JFS. The advanced FSS model would not only improve the overall segmentation performance, but also increase the likelihood of correctly identifying successful refinements. These approaches can substantially improve the success rate of JFS. As advancements in AI continue, particularly in the realm of few-shot segmentation models, there is still room for improving our framework with newly developed segmentation models. The continuous evolution of these models will enable more nuanced and precise segmentation, which can be leveraged within the JFS framework to achieve superior results.
In addition, there is also an opportunity to incorporate language descriptions as a novel form of guidance in evaluating refinement success as vision-language models (VLM) become increasingly sophisticated. Our framework has the potential to integrate semantic understanding into the refinement process by harnessing the descriptive capabilities of VLM, allowing more context-aware and accurate assessments. This potential integration of linguistic and visual information could lead to a more holistic and effective segmentation refinement strategy, extending the boundaries of what can be achieved with JFS. Generative AI is also a promising tool that can support the JFS framework by generating similar images and masks based on the query image, which can then be used as support masks. The aforementioned approaches would enable JFS to achieve higher accuracy in detecting even fine-detail differences.
Conclusion
We proposed JFS, a novel method for assessing the effectiveness of the segmentation refinement process. In this paper, we demonstrate that JFS has the capability to distinguish between successful and failed cases in SEPL segmentation refinement. The identification of successful refinements can be applied in various scenarios, leading to improved segmentation quality. Furthermore, JFS introduces a fresh perspective in the domains of few-shot segmentation and segmentation refinement by repurposing FSS for a unique application. However, we recognize that JFS underperforms in certain instances, indicating the need for further development. Future research could achieve better results by fine-tuning FSS models or adapting them specifically for the JFS framework. Moreover, exploring the integration of additional modalities, such as linguistic information, or leveraging generative AI could provide new avenues to enhance the JFS method. This continued exploration and refinement of JFS will contribute to expanding the boundaries of segmentation refinement and enriching its practical applications.